OpenStudy (anonymous):

write an equation of the line containing the given point. (0,8), 7x+4y=9

OpenStudy (owlcoffee):

There are infinite lines that can go through a point. That we can call "line family". You can take any other point and that point, and create a line that will go through the point you chose and the given point.

OpenStudy (anonymous):

ok so how would i find the equstion though

OpenStudy (owlcoffee):

I'll give you one of the many. I'll take "any point", wich means i'll chose the point "G" wich has coordinates (x,y). So no I have defined a point, I can use both to write a line equation, but let's begin by finding the slope. \[m=\frac{ y_g-8 }{ x_g-0 }\] \[m=\frac{ y_g-8 }{ x_g }\] So, now that we know the slope, we can use the equation that has this form: \[(y-y_p)=m(x-x_p)\] yp and xp are the coordinates of any point. but it has to be used by the slope, else this won't work. I'll chose the point (0,8). so: \[(y-8)=m(x-0)\] x-0 = x , And we have found the slope of the line we are looking for, so let's replace: \[(y-8)=(\frac{ y_g-8 }{ x_g })(x)\] So the end result is found by simplifying and putting the equation in a y=mx+b form. \[y-8=\frac{ (y_g-8)x }{ x_g }\] So then: \[y=\frac{ (y_g - 8 )x }{ x_g }+8\] finally: \[y=\frac{ (y_g-8)x+8x_g }{ x_g }\] That must represent the lequation of the line that goes through point (0,8) and any other point. Because Xg ang Yg are any point in the plane.